An urn contains 6 blue or white balls , 4 of them are blue , what is the pro sloth of getting white ball ? Please explain how you get the answer
A. 0.7
B. 1/3
C. 3
D. 1
Answer #1

Solution

Total number of balls n(T) = 6

Number of blue balls = n(B)=4

Number of white balls = n(W) = 2

P(selecting white ball ) = n(W)/n(T)

= 2/6 = 1/3

Hence answer is B. 1/3